The sum of the ages of Steve and Jeff is 69.In seven years, Steve age will be four less than twice Jeff's age . How old is each now?
Answer by ankor@dixie-net.com(22740)

s = Steve's present age
j = Jeff's
:
Write an equation for each statement
:
The sum of the ages of Steve and Jeff is 69.
s + j = 69
j = -s + 69
In seven years, Steve age will be four less than twice Jeff's age.
(s+7) = 2(j+7) - 4
s + 7 = 2j + 14 - 4
s = 2j + 10 - 7
s = 2j + 3
replace j with (-s+69) from the 1st statement
s = 2(-s+69) + 3
s = -2s + 138 + 3
s + 2s = 141
3s = 141
s = 141/3
s = 47 yrs is Steve's age
then
j = -47 + 69
j = 22 yrs is Jeff's age
:
:
Check solutions in the statement
"In seven years, Steve age will be four less than twice Jeff's age"
47 + 7 = 2(22+7) - 4
54 = 2(29)
54 = 58 - 4